OpenStudy (anonymous):

can someone help me factor 4x^2+20x-24?

OpenStudy (angela210793):

=4(x^2+5x-6)

OpenStudy (anonymous):

4(x+3)(x+2)

OpenStudy (amistre64):

step1; factor out the commons stepB; factor more stepZeta; present finished form :)

OpenStudy (amistre64):

the work involved can be minimal or maximal depending on the route you want to take

OpenStudy (amistre64):

but the answer doesnt really care how you get it :)

OpenStudy (amistre64):

raheen, i assume, typoed his reply since this is NOT the answer: 4(x+3)(x+2)

OpenStudy (anonymous):

\[x=-20+-\sqrt{(20)^2-4*4*(-24)}/2*4\] =[-20+-\sqrt{400+378}\]/8 \[\left[ -20+28\right]\]/8 \[8/8\] 1 ans using sri dhar acharya formula